OpenStudy (anonymous):

A training field is formed by joining a rectangle and two semicircles, as shown below. The rectangle is long and wide. Find the area of the training field. Use the value for , and do not round your answer.

OpenStudy (jmark):

Let the question is, A training field is formed by joining a rectangle and two semicircles, as shown below. The rectangle is 95m long and 68m wide. Find the area of the training field. Use the value 3.14 for PIE , and do not round your answer. Be sure to include the correct unit in your answer. solution: If the field is 68 m wide, then each semicircle has a diameter of 68 m , or a radius of 68/2 = 34 m. Area of two semicircles = Pi*R^2 = Pi(34)^2 = 3629.84 m^2 Area of Rectangle = LW = (95)(68) = 6460 m^2 Total Area = 6460 + 3629.84 = 10089.84 m^2 CBSE Sample Papers for Class 9: http://cbse.edurite.com/cbse-sample-papers/cbse-sample-papers-for-class-9.html